5) How many people are unemployed if the employment ratio is 75%, there are 90 million people 5) employed, and there are 20 million people not in the labor force? A) 5 million B) 15 million C) 10 million D) 20 million E) 0 millionshow work
Answer #1

ANSWER:

Employment ratio = no of employed people / total population

75% = 90,000,000 / total population

total population = 90,000,000 / 75%

total population = 90,000,000 / .75

total population = 120,000,000

total population = no of employed + no of unemployed + people not in labor force

120,000,000 = 90,000,000 + no of unemployed + 20,000,000

120,000,000 = 110,000,000 + no of unemployed

no of unemployed = 120,000,000 -110,000,000

no of unemployed = 10,000,000

so the correct answer is option c.